Please help me solve this problem.
Solve for x.
4x(x-4)-5x(x-3)=0
thank you!

Answer by sonya(5)

4x^2-5x^2-16x-15x=0
-x^2-x=0
-1(x^2+x)=0
x^2+x=0
-1^2+-1=0
x=-1